OpenStudy (anonymous):

A Librarian expansion begun in 2002 was expected to cost 107mill. the cost ended up being 44 mill over the budget. by what percent was the actual cost exceeded from the estimate

OpenStudy (anonymous):

107*percentage = 107 + 44 percentage = (107 + 44)/107 = 1.411 1.411* 100 = 141.1 % , 41.1% over